OpenStudy (anonymous):

rationalize the denominator of 7/(3 + √3)

jimthompson5910 (jim_thompson5910):

Multiply top and bottom by 3-√3 to get \[\Large \frac{7(3-\sqrt{3})}{(3+\sqrt{3})(3-\sqrt{3})}\] \[\Large \frac{21-7\sqrt{3}}{3^2-(\sqrt{3})^2}\] \[\Large \frac{21-7\sqrt{3}}{9-3}\] \[\Large \frac{21-7\sqrt{3}}{6}\] So \[\Large \frac{7}{3+\sqrt{3}} = \frac{21-7\sqrt{3}}{6}\]

OpenStudy (anonymous):

21-7 square root 3/ 6..